OpenStudy (anonymous):

one more please help with proving this identity: (1/tan2x)+(cos2x/cot2x)=cot2x+sin2x

OpenStudy (anonymous):

Notice that 1/tan2x = cot2x. So re-write this as:\[L.S=\cot(2x)+\frac{ \cos(2x) }{ \cot(2x) }=\cot(2x)+\sin(2x)=R.S\]@samancha85

OpenStudy (anonymous):

@samancha85

OpenStudy (anonymous):

ok